MalayaJanae:

Jimmy made 16 quarts of lemon juice. How many liters of lemon juice did Jimmy make? (5 points) 1 gallon = 4 quarts 1 gallon = 3.785 liters

MalayaJanae:

Please Explain Before Giving Out Answers!!!!

supie:

To convert quarts to liters I believe you have to divide the amount of quarts by 1.057

supie:

@jhonyy9

jhonyy9:

use the rule of 3 simple and write in this way 1 gal. = 4 quarts x gal = 16 quarts cross multiplie and calcule x x = 16*1/4 = 16/4 = 4 gal 1 gal =3,785 liters 4 gal = x lit. cross multiplie and calcule x x = 4*3,785/1 = 4*3,785 = ? liters hope helped easy understandably @supie

draebrewer:

what do u think the answer is

draebrewer:

1 gallon = 4 quarts 1 gallon = 3.785 liters 16 quarts = 4 gallons 4 x 3.785 =

draebrewer:

i worked it out for you, but i wont tell u the answer yet